### v1

# VALUE: Value-Aware Large Language Model for Query Rewriting via Weighted Trie in Sponsored Search

Abstract.
In the realm of sponsored search advertising, matching advertisements with the search intent of a user’s query is crucial. Query-to-bidwords(i.e. bidding keywords) rewriting, which involves transforming user queries into keywords for bidding, is a vital technique that has garnered significant attention from both industry and academia. Recently, with the prevalence of large language models (LLMs), generative retrieval methods have proven effective in producing high-relevance rewrites. However, we have identified a significant limitation in existing approaches: While fine-tuning LLMs for specific domains enhances semantic relevance, these models have no perception of the intrinsic value of their generated outputs, such as commercial value. Therefore, after supervised fine-tuning (SFT), a reinforcement learning from human feedback (RLHF) phase is often employed to address this issue. Nevertheless, traditional preference alignment methods often face challenges in aligning fine-grained values and are susceptible to overfitting, which diminishes the effectiveness and quality of the generated results. To address these challenges, we propose VALUE (Value-Aware Large language model for qUery rewriting via wEighted trie), the first framework that ensures the generation of high-value and highly relevant bidwords. Our approach utilizes weighted trie, an innovative modification of the traditional trie data structure. By modulating the LLM’s output probability distribution with value information from the trie during decoding process, we constrain the generation space and guide the trajectory of text production. Our method not only addresses fine-grained value alignment but also effectively reduces the hallucination issues often encountered with LLMs. Offline experiments demonstrate the effectiveness of our method in semantic matching and preference alignment, showing a remarkable improvement in the value attribute by more than fivefold. Online A/B tests further revealed that our Revenue Per Mille (RPM) metric increased by 1.64%. VALUE has been deployed on our advertising system since October 2024 and served the Double Eleven promotions, the biggest shopping carnival in China.
Query Rewrite; Semantic Matching; Generative Retrieval; Large Language Model;
These authors contributed equally to this work and should be considered co-first authors.
This work was done when the first author was an intern at Alimama.
copyright: acmlicensedjournalyear: 2024conference: Make sure to enter the correct conference title from your rights confirmation emai; TBD; TBDisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Query reformulationccs: Information systems Top-k retrieval in databasesccs: Computing methodologies Natural language processing
1. Introduction
Figure 1. System Overview of our proposed VALUE Framework includes three parts: ”SFT and Alignment”, ”Weighted Trie Construction and Weight Momentum Update”, and ”Collaborative Inference”. During the LLM decoding process, we derive a value distribution from a weighted trie to constrain the generation space, and subsequently merge this with another distribution to produce the final output distribution.
[…]
Sponsored search is a crucial element of modern search engines, where advertisers bid on relevant bidwords to display their ads alongside original search results. Upon receiving user’s query, it is essential to match the query with relevant bidwords (Fain and Pedersen, 2006), and subsequently identify suitable ads associated with those bidwords. This process is a critical component in the ad retrieval pipeline, driving the major revenue of search engines (Jansen and Mullen, 2008). The success of sponsored search depends on the search engine’s ability to accurately identify bidwords that closely align with user intent. However, users often face challenges when formulating queries, leading to suboptimal search experiences. These challenges are intensified when users do not use precise or correct terminology, employ synonyms, or mix languages in their search phrases due to varying levels of language proficiency. Additionally, search terms might be misspelled or overly general, complicating the retrieval of relevant ads. For example, a user might search for ”spring fashion,” which is broad and could correspond to bidwords like ”2024 spring collection” or ”women’s spring trends”. Each query reflects different intents but lacks clarity without additional context. As e-commerce platforms continue to grow in both scale and diversity, ensuring accurate and relevant ad retrieval becomes increasingly challenging, necessitating the need for advanced query rewriting and bidword matching techniques to better align user queries with advertiser bidwords.
In query rewriting, two primary paradigms exist: discriminative methods, which reformulate queries by the retrieval of similar terms and generative methods, which utilize language models for direct query-to-bidword transformation. Discriminative methods leverage sparse and dense retrieval techniques to find relevant bidwords. Sparse retrieval techniques, such as BM25 (Robertson et al., 2009), employ exact lexical matching through high-dimensional term-frequency vectors. Their reliance on exact lexical matching leads to semantic brittleness. This rigidity prevents them from recognizing semantic equivalences between different expressions, resulting in significant retrieval failures when faced with semantically similar but lexically divergent queries. Dense retrieval approaches overcome the limitations of sparse methods by utilizing low-dimensional semantic embeddings generated by neural encoders like sBERT (Reimers, 2019). Dense retrieval, while effective in capturing semantic relationships, suffers from significant limitations, particularly the Matthew Effect and long-tail challenges. High-frequency queries dominate model training due to abundant click logs, leading to over-optimized embeddings for head queries while neglecting tail queries with sparse data. Furthermore, dense retrieval models struggle with distribution shifts and require costly strategies like meta-learning or data augmentation to address long-tail issues (Peng et al., 2023).
Recently, the rise of generative retrieval has garnered widespread attention. Wang (Wang et al., 2024) utilizes LLMs for one-step query-to-keyword rewriting, using a trie during the decoding process. Despite improvements over traditional information retrieval (IR) baselines, existing generative approaches using LLMs for query-to-bidword rewriting tasks have notable limitations. A key limitation is their inability to recognize the value of the generated results, which, in our scenario, pertains to the outcomes they can deliver. Models trained solely through supervised fine-tuning on online logs lack awareness of the value of the results. Due to the training methodology and the SFT loss function design, the model treats each query-bidword pair in the training samples equally, failing to learn any preference or value information. Although there is a great deal of work on Reinforcement Learning from Human Feedback (RLHF), such as PPO, DPO, KTO, and ORPO(Ouyang et al., 2022; Rafailov et al., 2024; Ethayarajh et al., 2024; Hong et al., 2024), which can partially mitigate the value perception issues of SFT models, these methods also have significant drawbacks. Firstly, they require substantial computational resources for training. However, the effective Cost Per Mille (eCPM) value of bidwords is a statistical measure that frequently fluctuates. When these fluctuations occur, the previous alignment efforts become obsolete, and the model needs to be retrained. Secondly, these training methods are prone to overfitting (Meng et al., 2024), especially in our scenario. Unlike traditional alignment objectives for LLMs, such as harmlessness and usefulness, which can be achieved using a small dataset, the value of a bidword is completely independent of its textual content. Therefore, the model must be trained individually for each bidword. Our experiment on DPO show that models trained in this manner tend to overfitting. There is an urgent need for a new paradigm to address the challenges above.
Through our observations, we have identified an opportunity in the sponsored search advertising scenario to enhance the value of bidwords without compromising their relevance. The value of a bidword does not necessarily correlate with its literal meaning. Two bidwords might differ by only a symbol or character, yet exhibit substantial differences in value. Therefore, there is an opportunity to increase the value of the top $K$ bidwords while minimally affecting relevance by prioritizing those with the same level of relevance but higher value over others. Thereby, we present VALUE, an innovative reward-guided inference framework that integrates value-awareness into the decoding process via a novel weighted trie structure. We design algorithms for the construction and update of the value of trie node. During decoding process, VALUE obtain reward information from for the next tokens the weighted trie. This information is used to adjust the output distribution, thereby sampling the subsequent token. By this reward-guided search, the framework aims to maximize the reward of the outputs. We used a multi-task fine-tuning strategy to equip the model with essential query rewriting skills. To achieve better value alignment, we designed an optional post-training alignment phase. To evaluate the effectiveness of our proposed method, we conducted offline experiments using historical online logs. Compared to the current online service model, our method demonstrated a 9.3% improvement in relevance. Additionally, online A/B tests showed increases in key profit metrics, with Cost rising by 1.53% and Revenue Per Mille (RPM) increasing by 1.64%. Given these significant improvements, our method has been fully deployed in the online service. Our contributions can be summarized as follows:
  1. (1)
    Pioneering Query Rewriting Framework via Collaborative Inference: We present a groundbreaking query rewriting framework that integrates LLMs with weighted trie, uniquely designed to simultaneously optimize for both high value and semantic relevance. This framework represents the first of its kind to effectively balance these two critical aspects.
  2. (2)
    Renovation of Trie: We have reinvented the trie data structure, by seamlessly integrating value information directly into its nodes. This advancement is further supported by our proposed algorithms for trie construction and momentum update of node value.
  3. (3)
    Post-Training Alignment for Fine-Grained Reward Optimization: We propose an post-training alignment algorithm that fine-tunes the model to better align with fine-grained reward attributes.
2. Related Works
2.1. Query Rewriting
The process of query rewriting, often termed query expansion or reformulation, is essential in enhancing e-commerce search technologies. It profoundly affects the shopping experience of users and the financial performance of e-commerce platforms. This approach can be generally classified into two classes: discriminative methods and generative methods.
Discriminative methods approach query rewriting as a retrieval task. Pseudo-relevance feedback methods (Xu and Croft, 2017b; Cao et al., 2008; Tao and Zhai, 2006; Xu and Croft, 2017a) enhance queries by identifying expansion terms from the top-ranked documents of an initial query, combining global corpus analysis with local feedback. Although these methods effectively address word mismatches, they are susceptible to semantic drift due to noisy or irrelevant top results. To address these challenges, researchers suggest utilizing a well-constructed thesaurus as a candidate set for query rewriting. (Bhogal et al., 2007; Mandal et al., 2019) However, they caution that the effectiveness of these methods is highly dependent on the thesaurus’s quality. If the thesaurus is inadequate, it may lead to query semantic drift, where the intended meaning of the query is compromised. Additionally, some approaches (Antonellis et al., 2008; Cui et al., 2002; Li et al., 2022) focus on generating candidate rewrites by leveraging search logs, using similar terms from users’ search histories as extensions. However, these search logs inherently display a bias towards popular queries due to the Matthew effect. Consequently, the training data derived from this method may fall short in optimizing for less frequently searched long-tail queries.
Generative methods, on the other hand, focus on the direct transformation of queries into keywords (Wang et al., 2024; Lee et al., 2018; Qi et al., 2020; Mohankumar et al., 2023). These methods involve transforming the input query directly into multiple bidwords, utilizing auto-regressive language models combined with trie-based decoding techniques to restrict the output space during inference. However, while previous methods ensure semantic relevance between input and output, they fall short in addressing value—a critical attribute in sponsored search advertising—which remains an open challenge. Our approach, as detailed in the following sections, effectively tackles this challenge by introducing innovative techniques that not only ensure semantic relevance but also optimize for value. By addressing this critical gap, our research has the potential to revolutionize sponsored search advertising, leading to more effective ad targeting and increased user satisfaction.
2.2. Preference Alignment
Reinforcement Learning from Human Feedback (RLHF) represents a pivotal technique for aligning LLMs with human preferences and values (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022), commonly applying the Bradley-Terry model (Bradley and Terry, 1952) to estimate the probability of a pairwise competition between two independently evaluated instances. The conventional RLHF pipeline is structured into three distinct phases: supervised fine-tuning (Zhou et al., 2024; Taori et al., 2023; Conover et al., 2023; Ding et al., 2023), reward model training (Gao et al., 2023; Chen et al., 2024; Havrilla et al., 2024b; Leike et al., 2018), and policy optimization (Schulman et al., 2017; Anthony et al., 2017). Notably, Proximal Policy Optimization (PPO) (Schulman et al., 2017) is extensively employed during the policy optimization phase. The RLHF framework has found broad applicability across various domains, including the mitigation of toxicity (Amini et al., 2024; Korbak et al., 2023), the assurance of safety (Dai et al., 2023), the enhancement of helpfulness (Wei et al., 2024) and the improvement of model reasoning capabilities (Havrilla et al., 2024a). Recent studies (Casper et al., 2023) have underscored the challenges inherent in the entire RLHF pipeline, spanning from the collection of preference data to the training of models. Furthermore, subsequent research has illuminated potential biases introduced by RLHF, such as the generation of verbose outputs (Dubois et al., 2024; Singhal et al., 2023; Wang et al., 2023).
3. Methodology
Traditional methods, such as standard beam search or top-k sampling, prioritize candidates with higher probabilities, often neglecting their value attributes. Although these approaches are effective in numerous contexts, they may fall short in applications where the value attributes are important. VALUE addresses this limitation by incorporating a Weighted Trie that allows for value-aware token selection. At each step of the decoding process, VALUE modifies the model’s output probabilities based on the value obtained from the trie, prioritizing tokens of higher value among those already highly relevant.
Figure 2. Collaborative Inference of VALUE framework. When generating the next token, we obtain two output distribution from the LLM and weighted trie. We exclude tokens not present in the value distribution and adjust the probabilities of the remaining tokens. Then, we sample the next token from the modified distribution.
3.1. System Overview
As can be seen from Figure 1, our entire system comprises the following key modules: ”LLM SFT and Alignment,” ”Weighted Trie Construction and Weight Momentum Update,” and ”Collaborative Inference.” For each generated token, we obtain the probability distribution from the LLM. Concurrently, we retrieve all possible candidates under the current prefix from the weighted trie along with their respective values. By integrating these values with the probability distribution, we derive a new probability distribution and sample the next token from it.
3.2. Weighted Trie
3.2.1. Definition and Construction of Weighted Trie
Our Weighted Trie is an renovation of the traditional trie data structure. Each node in the Weighted Trie, referred to as a WeightedTrieNode, contains additional attributes to store weight-related information. The primary attributes of a WeightedTrieNode are:
  • children: A dictionary where the keys are integers representing tokens, and the values are child nodes of type WeightedTrieNode.
  • mean: A floating-point number representing the average weight of the node’s children.
  • max: A floating-point number representing the maximum weight among the node’s children.
  • is_word: A boolean value indicating whether the node represents a complete word.
Construction of Weighted Trie: the construction of the Weighted Trie involves initializing the root node, inserting bidwords with their corresponding values, and updating the weights of the nodes. The pseudocode in Algorithm 1 demonstrates the construction process.
Algorithm 1 Construct Weighted Trie
[…]
  Output: Weighted Trie
[…]
     end for
[…]
  end for
  Update weights in the weighted trie using post-order traversal:
  for each node in post-order do
[…]
  end for
3.2.2. Weighted Trie Momentum Update
Momentum updates are essential due to the frequent changes in eCPM and the evolving bidword space. The process is delineated as follows:
Tokenization and Node Insertion:
Each bidword undergoes tokenization, converting it into a sequence of token IDs. These IDs facilitate the insertion or updating of nodes within the trie. The update process for a node is governed by:
(1) $V_{\text{new}}=\max(eCPM_{\text{new}},V_{\text{old}})$
(2) $W_{\text{new}}=\alpha\cdot eCPM_{\text{new}}\ +\ \beta\cdot W_{\text{old}}$
where $\alpha$ and $\beta$ are hyperparameters that control the update rate, with the constraint $\alpha+\beta=1$. V is the max attribute of the node and W is the mean attribute of the node. If the node is absent, it is initialized as:
(3) $W_{\text{new}}=eCPM_{\text{new}},\ V_{\text{new}}=eCPM_{\text{new}}$
Bottom-up Update:
Following node insertion, a bottom-up iteration is executed to update all intermediate nodes, ensuring that the entire trie structure accurately reflects the most recent eCPM values.
This method enables the trie to dynamically adjust to variations in the bidword space and eCPM values over time, preserving the framework’s efficacy without necessitating frequent retraining, maintaining stability and performance in a dynamic e-commerce environment.
3.3. Collaborative Inference
As shown in Figure 2, when generating the $t$-th token, we insert the query into the prompt template and concatenate it with the bidword prefix to form the input of LLM to get the output probabilities. We use generated bidword prefix to retrieve all child nodes from the trie. The value of these child nodes is computed using the following formula:
(4) $\mathcal{V}_{k}=\alpha\cdot\text{mean}_{k}+\beta\cdot\text{max}_{k}$
Here, $\alpha$ and $\beta$ are tunable hyperparameters, both set to 0.5, allowing us to prioritize paths with higher average values or to explore bidwords with the maximum potential value, depending on our specific needs.
We then apply the softmax function to these values to obtain the normalized value $\hat{\mathcal{V}}_{k}$. Next, we mask out all probabilities in the LLM-generated distribution that are not children of the current node. For the remaining probabilities, we adjust them by weighting with the normalized value $\hat{\mathcal{V}}_{k}$ and a depth-dependent factor $\theta$. The adjusted probability for each token $k$ is given by:
(5) $p(k\mid x_{<t})=p_{\text{LLM}}(k\mid x_{<t})\cdot(1+\mathcal{V}_{k}\cdot\theta)$
where $\theta$ is designed to vary with the depth of the trie, playing a crucial role in balancing relevance and value in our model’s output. At shallower levels of the trie, $\theta$ is set to smaller values to minimize the influence of value on the output, thereby preventing the model from favoring high-value but less relevant options. As we delve deeper into the trie, $\theta$ increases, allowing us to prioritize higher-value candidates from a pool of already highly relevant choices. Consequently, we derive a modified probability distribution $p(k\mid x_{<t})$, from which the next token is sampled as usual.
3.4. Multi-task Fine-Tuning
As no open-source LLMs are trained to perform query rewriting for the e-commerce scenario, they lack an awareness of the bidword space. Consequently, they are unable to effectively generate bidwords. To address this, we need to perform supervised fine-tuning (SFT) on these models to inject domain-specific knowledge into the model and enhancing their capability to understand and effectively rewrite e-commerce queries for improved user experience and search accuracy.
Query Rewriting Dataset: To construct our training dataset, we initially obtain query-bidword pairs from online logs. Specifically, when a user searches for a query $x$ on our e-commerce platform, the system logs the corresponding bidwords $Y=\{y^{1},y^{2},\ldots,y^{n}\}$ that are associated with the advertisements the user interacts with. From this list, we select the top-ranked bidword $y^{1}$ as the gold standard candidate to construct our initial dataset $D$ with $N$ samples:
(6) $D=\{(x_{i},y_{i}^{1})\}_{i=1}^{N}\quad\text{where}\quad x_{i}\sim p(x)$
where $p(x)$ denotes the query distribution in our search engine.
We apply a relevance filter to $D$ using a representation model fine-tuned from an open-source embedding model. The relevance score $\text{rel}(x,y)$ is computed as the cosine similarity between the embeddings of $x$ and $y$:
(7) $\text{rel}(x,y)=\frac{\text{emb}(x)\cdot\text{emb}(y)}{\|\text{emb}(x)\|\|\text{emb}(y)\|}$
We then filter the dataset $D$ to obtain $D_{r}$ ensuring that the relevance score exceeds a threshold $\tau_{\text{rel}}$:
(8) $D_{r}=\{(x_{i},y_{i})\mid(x_{i},y_{i})\in D,\,\text{rel}(x_{i},y_{i})>\tau_{\text{rel}}\}$
For head queries, the number of corresponding bidwords can be in the tens of thousands. To ensure class balance, we truncate the bidwords for each query based on their value.
Fine-tuning Tasks: We employ two types of supervised fine-tuning (SFT) tasks to enhance the model’s understanding and performance in query rewriting for e-commerce.
Task 1: <prompt + query + bidword>
This task aims to enhance the model’s spatial awareness of the bidword space and the association between queries and bidwords. Subsequently, the model will develop fundamental rewriting skills.
Task 2: <prompt + query + bidword_list>
In this task, the model is trained with lists of bidwords sorted in descending order based on their eCPM values. The bidword_list for each query is structured as follows:
[…]
3.5. LLM Post-training Alignment with Weighted Trie
Although conventional RLHF approaches can effectively integrate with our framework, we propose a more tailored solution. Our Weighted DPO (WDPO) approach directly incorporates the economic value of bidwords into the alignment process, leveraging the eCPM values for bidwords. It is important to note that WDPO is optional as the performance improvements it offers are not as significant relative to computational cost when compared to SFT and the varying nature of bidwords. Therefore, WDPO can be considered when resources permit, providing an additional layer of refinement to the training process.
The alignment process proceeds as follows:
  1. (1)
    Identify relevant bidwords from online logs and calculate their corresponding eCPM values for each query in our dataset.
  2. (2)
    Randomly sample two bidwords $(y_{w},y_{l})$ with differing eCPM values, ensuring:
    (9) $|eCPM(y_{w})-eCPM(y_{l})|>\tau$
    where $\tau$ is a predefined threshold, $y_{w}$ represents the bidword with higher eCPM, and $y_{l}$ the lower.
  3. (3)
    Process the query and selected bidwords through both policy and reference models to obtain probabilities: $\pi_{\theta}(y_{w}|x)$, $\pi_{\theta}(y_{l}|x)$, $\pi_{ref}(y_{w}|x)$, and $\pi_{ref}(y_{l}|x)$.
  4. (4)
    Apply the modified DPO loss function:
    (10) $\mathcal{L}_{WDPO}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\bigg{[}$ $\log\sigma\bigg{(}\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}$
    $-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)}\bigg{)}\cdot w\bigg{]}$
    where $w$ is a weight function based on KL divergence:
    (11) $w=\exp(-KL([\pi_{\theta}(y_{w}|x),\pi_{\theta}(y_{l}|x)]\parallel[P(y_{w}),P(y_{l})]))$
    and $P(y_{i})$ represents normalized eCPM values:
    (12) $P(y_{i})=\frac{eCPM(y_{i})}{\sum_{j\in\{w,l\}}eCPM(y_{j})}$
3.6. Online Depolyment
In the context of deploying LLMs for online serving, the balance between computational efficiency and latency is paramount. Our approach leverages the Qwen2-7B model for offline inference, targeting head queries that constitute 50% of the page views (PV) in our system. This offline process generates the top 500 results per query, which are subsequently cached for rapid access. For mid-tail and long-tail queries that do not benefit from the cached results, we employ the Qwen2-1.5B model in an online serving capacity. This model is optimized to deliver a latency of 50 milliseconds, efficiently producing the top 50 results per query. In the future, we plan to integrate user behavior sequence data into our online serving framework to enhance personalized recommendation systems and gradually remove the offline part.
4. Experiments
4.1. Datasets
[…]
Test Dataset: Our test dataset was constructed by randomly sampling 30,000 queries from online logs. These queries were stratified into three categories: 40% head queries, 40% torso queries, and 20% tail queries.
4.2. Evaluation Metrics
Offline Metrics: To comprehensively evaluate our query rewrite approach, we utilize several offline metrics:
  1. (1)
    Hit Rate: A precision-oriented metric measuring the alignment between generated rewrites and user-clicked bidwords. For a query set $\mathcal{Q}$, it is defined as:
[…]
    where $\mathcal{R}_{q}^{K}$ represents the top-$K$ rewrite candidates for query $q$, and $\mathcal{C}_{q}$ denotes the set of clicked bidwords for $q$. The numerator counts total hits across all queries, while the denominator normalizes by the total clickable bidwords.
  2. (2)
    Relevance Score: A semantic alignment measure between rewritten queries and the original query intent, typically quantified via pretrained language model similarity (e.g., BERTScore).
  3. (3)
    Spearman Rank Correlation ($\rho$): A non-parametric measure of rank consistency between model-generated and ground-truth bidword rankings. For $n$ observations:
    (14) $\rho=1-\frac{6\sum_{i=1}^{n}d_{i}^{2}}{n(n^{2}-1)}$
    where $d_{i}$ is the rank difference for the $i$-th bidword pair. The coefficient $\rho\in[-1,1]$ reflects the monotonicity of the predicted rankings.
Online Metrics: To evaluate the model’s performance, we use three key metrics: cost, which is the total advertiser expenditure for clicks within a specific traffic segment, revenue per mille (RPM), which measures the revenue generated per thousand ad impressions and Page View Relevance (PV rele),
4.3. Implementation Details
Multi-task Fine-tuning: We fine-tuned the Qwen2-7B model using a learning rate of 1e-5 and a cosine learning rate scheduler. The process was optimized with AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and a weight decay of 0.001. Gradient accumulation was set to 4. We used a batch size of 128 and applied the DeepSpeed ZeRO Stage 2 parallelism strategy with bfloat16 precision. The training was performed on 32 NVIDIA H20 GPUs for a total of 12 hours.
Post-training Alignment: For the post-training alignment phase, we use our fine-tuned Qwen2-7B model. The alignment phrase was carried out with a learning rate of 1e-6 and a batch size of 64. We set the $\beta$ parameter to 0.1. The training was executed on 16 NVIDIA H20 GPUs for a duration of 3 hours.
4.4. Offline Experiments
Table 1. Comparative Performance Analysis of Various Models. The eCPM metric is used as an approximate estimation of reward, serving as a reference for the effectiveness of weighted trie collaborative inference.
Model hitrate@50 hitrate@500 Spearman’s $\rho$ Relevance OOVR eCPM
OnlineModel 59.93% 85.41% 0.02 68.15% 0% 8676
DeepNB 44.16% 79.82% 0.13 46.26% 0% 12852
BEQUE 39.26% 62.85% 0.08 74.37% 38.41% 7582
raw Qwen2-7B 12.88% 18.37% -0.03 59.81% 68.63% 2937
SFT Qwen2-7B 38.18% 63.71% 0.02 76.33% 34.85% 7439
SFT+Trie 59.11% 86.38% -0.01 75.31% 0% 7270
SFT+Trie+DPO 38.61% 62.19% 0.05 69.48% 0% 28391
SFT+WeightedTrie(VALUE) 60.37% 91.52% 0.46 74.55% 0% 59803
SFT+WeightedTrie+DPO 50.35% 77.73% 0.23 70.18% 0% 38301
SFT+WeightedTrie+WDPO(VALUE+WDPO) 62.67% 91.91% 0.56 73.49% 0% 63775
In this section, we present the results of offline experiments conducted to evaluate the performance of our proposed models against several baseline models. The comparative performance analysis is summarized in Table 1. We compared our models with the following baselines: first, the Online Model, which represents the current state-of-the-art multi-channel recall system deployed online; second, the DeepNB model (Chen et al., 2020), which employs a vector-based approach focusing solely on value; and finally, the BEQUE model (Peng et al., 2023), which utilizes a LLM fine-tuned with multi-task SFT and PRO alignment, representing the latest query rewriting approach. The performance metrics used for evaluation include hitrate, eCPM, Spearman’s $\rho$, Relevance, and OOVR(Out-Of-Vocabulary Rate. To a certain extent, it is also equivalent to the hallucination rate of LLMs.). These metrics provide a comprehensive assessment of the models’ effectiveness in terms of query rewriting, relevance, and OOVR.
4.4.1. Comprehensive Evaluation of the VALUE Framework
The VALUE framework we propose is designed to optimize online RPM revenue by utilizing offline eCPM values as a reference for learning effectiveness, rather than as direct revenue indicators. Our subsequent online experiments have demonstrated the validity and rationality of this modeling approach.The results presented in Table 1 clearly illustrate the significant effectiveness of our VALUE framework. Notably, the eCPM metric shows a remarkable improvement, serving as a strong indicator of offline learning success. Compared to the model trained with Supervised Fine-Tuning (SFT) on the same dataset, our VALUE model achieves a 722.6% increase in eCPM. Furthermore, when compared to the current online service model, our framework exhibits a 589.3% enhancement in eCPM. Importantly, the relevance metric for our VALUE model does not decline when compared to the SFT model; in fact, it demonstrates an improvement of 6.4 percentage points over the online service model. This observation substantiates our earlier hypothesis that relevance and value are decoupled in this domain, allowing us to reorder the generation sequence of bidwords to enhance both metrics simultaneously without trade-offs.
Compared to the DeepNB model, which focuses on efficiency-oriented rewriting using click data to maximize value, our VALUE framework demonstrates superior performance across multiple dimensions. While the DeepNB model effectively maximizes value, it suffers from poor relevance metrics. In contrast, our VALUE framework not only maintains high relevance but also significantly outperforms DeepNB in terms of value attributes. Specifically, the eCPM metric for our VALUE model is substantially higher than that of DeepNB. When compared to the BEQUE model, our framework achieves comparable relevance metrics but excels in generating high-value bidwords. This is evidenced by the eCPM metric, where VALUE achieves a value of 59803, far surpassing BEQUE’s 7582. Additionally, our VALUE framework demonstrates a superior hit rate, with hitrate@500 reaching 91.52%, compared to BEQUE’s 62.85%. This significant improvement in hit rate underscores the effectiveness of our approach in generating bidwords that are not only relevant but also highly likely to be clicked by users. It is noteworthy that the BEQUE method suffers from a high Out-Of-Vocabulary Rate (OOVR) of 38.41%, which adversely affects its hit rate.
Moreover, our VALUE framework outperforms the online service model across all evaluated metrics, further validating the efficacy of our approach. The comprehensive improvements in eCPM, relevance, and other metrics underscore the robustness and effectiveness of our method in generating high-value and contextually relevant bidwords for e-commerce query rewriting tasks.
Figure 3. Trade-off between RPM and PV rele through $\theta$.
[…]
Table 2. Comparative Performance Analysis of Hyperparameter.
Hyperparameter($\theta$) hitrate@500 Relevance
$\theta_{1}$ 86.38% 75.31%
$\theta_{2}$ 89.65% 75.04%
$\theta_{3}$ 91.52% 74.55%
$\theta_{4}$ 88.63% 69.85%
$\theta_{5}$ 73.71% 60.53%
4.4.2. Enhanced Performance in Fine-Grained Value Alignment
Compared to traditional alignment methods such as DPO and Preference Ranking Optimization (PRO) used in the BEQUE model, our VALUE framework demonstrates a significant lead in the Spearman rank correlation metric. This indicates that conventional value alignment methods struggle with fine-grained value alignment tasks. Specifically, the Spearman’s $\rho$ for our VALUE model is 0.46, which is substantially higher than the 0.05 achieved by DPO model and 0.08 by BEQUE.
Additionally, our framework seamlessly integrates with conventional model alignment methods. When we combined our proposed WDPO with the VALUE framework, we achieved the highest Spearman’s $\rho$ of 0.56.
DPO suffers from significant overfitting issues, which severely impact its performance in terms of hit rate and relevance. In our experiments, the SFT + DPO + Trie model exhibited a hitrate@500 of only 62.19% and a relevance score of 69.48%, both of which are considerably lower than those achieved by our VALUE framework. DPO faces inherent challenges when dealing with positive and negative samples that are lexically very similar. The overfitting problem in DPO arises because the method tends to excessively optimize for the specific preference pairs seen during training, leading to a decline in generalization capability. DPO enhances the model’s belief in specific events. Therefore, when we use DPO in conjunction with our VALUE framework, the improvement in eCPM is not significant. Specifically, the eCPM for the VALUE + DPO model is 38301, which is only marginally higher than the 28391 achieved by the SFT + DPO + WeightedTrie model.
Furthermore, the improvement in the Spearman rank correlation metric with DPO is not significant. The SFT + DPO + Trie model achieved a Spearman’s $\rho$ of only 0.05, which is marginally better than the baseline but far from satisfactory. This limited improvement underscores the inherent limitations of DPO in handling fine-grained value alignment tasks.
4.5. Exploration of Hyperparameter Configurations
In this section, we provide a concise analysis of the impact of the hyperparameter $\theta$ on system performance. As shown in Table 2, an increase in $\theta$ results in a decrease in relevance. Figure 3 visually represents the dynamics of this trade-off. Initially, optimizing for value leads to an increase in RPM; however, when relevance becomes significantly low, RPM ultimately decreases. This underscores the critical importance of meticulous hyperparameter tuning to ensure alignment with specific application objectives. For a detailed exploration of the hyperparameter $\theta$, please refer to Appendix A.
4.6. Scaling Laws in Model Performance
In our investigation of model performance across varying scales, we conducted experiments using models ranging from 0.5 billion to 65 billion parameters. Our findings consistently demonstrate that larger models exhibit superior performance across all evaluated metrics. Specifically, as model size increases, there is a notable improvement in hitrate, indicating enhanced capability in generating high-value and relevant bidwords. This trend aligns with the theoretical expectations of scaling laws, which suggest that larger models can capture more complex patterns and nuances in data, thereby improving overall effectiveness.
Table 3. Comparative Performance Analysis of Base Model Size.
Model Size hitrate@500 Relevance
0.5B 68.32% 61.85%
1.5B 87.49% 71.29%
7B 91.52% 74.55%
14B 92.69% 76.68%
65B 93.98% 77.99%
4.7. Online Experiments
To further validate the effectiveness of our proposed VALUE framework, we conducted 14-day online A/B tests on our system, focusing on cost and RPM. Results are summarized in Table 4 and Table 5.
Table 4. Online A/B Test Results
Query Type Cost RPM PV rele
VALUE all queries +1.53% +1.64% +0.32pt
VALUE head queries +0.97% +1.03% +0.10pt
VALUE torso queries +1.68% +1.74% +0.33pt
VALUE tail queries +2.35% +2.66% +0.75pt
The online A/B tests corroborate the offline findings, consistently demonstrating improvements in both cost and RPM across all query types. Our framework exhibits particularly notable enhancements for torso and tail queries, which are typically characterized by greater variability and complexity. This underscores the framework’s robust generalization capabilities. As illustrated in Table 5, our model excels in discerning the intent behind very long and complex queries. The observed improvements in RPM and relevance for medium to super long queries underscore the model’s proficiency in managing intricate and challenging query structures.
Table 5. Comparative Performance Analysis of Query Length Layer.
Query Length Cost RPM Relevance
Short Query +0.41% +0.56% +0.02pt
Medium Query +0.88% +0.98% +0.30pt
Long Query +1.26% +1.56% +0.79pt
Super Long Query +1.17% +1.55% +0.54pt
4.8. Case Study
Figure 4. Case Study of VALUE Framework
[…]
In this section, as illustrated in Figure 4, we examine a specific case involving the user query ”fatty’s drink.” Previously, without the use of LLMs, our system could only match the query at the keyword level, specifically ”drink.” With the integration of LLMs, we gained a deeper understanding of the user’s intent, identifying a preference for low-sugar, low-calorie beverages. Additionally, by employing a weighted trie, we ensured that the advertisements presented were not only closely aligned with the user’s intent but also of high value.
5. Conclusion
In this study, we introduced VALUE, an innovative framework designed to enhance query rewriting in sponsored search advertising by integrating value-awareness into the generation process. Our approach addresses the significant challenge of producing rewrites that are not only semantically relevant but also economically valuable—a gap that existing methods often fail to bridge. By embedding value information directly into the LLM’s decoding process through a novel weighted trie structure, we effectively steer generation towards high-value bidwords without substantially compromising relevance.
Comprehensive experiments underscore the effectiveness of VALUE in enhancing the economic attributes of bidword generation. Offline evaluations demonstrate a significant increase in eCPM, outperforming traditional methods while preserving high relevance scores. The improvement in Spearman rank correlation confirms the superior ranking quality achieved through our fine-grained value alignment. Furthermore, online A/B testing validates the real-world efficacy of VALUE, highlighting notable improvements in key performance indicators such as cost and RPM. Our framework’s versatility also suggests potential applications beyond sponsored search advertising, wherever the generation space is fixed and reward signals are present.
Appendix A Hyperparameter Details
In this appendix, we delve into the specifics of the hyperparameter $\theta_{d}$, which is designed to vary with the depth of the trie. The selection of $\theta_{d}$ is critical in modulating the influence of value relative to relevance as the trie depth changes. The configurations explored in our research are as follows:
  • $\theta_{1}$: [0, 0, 0, 0, 0, 0, 0, …], indicating that the weighted trie is ineffective and the original LLM reasoning result is used.
  • $\theta_{2}$: [1, 1, 1, 1, 1, 1, 1, …], implying equal weighting of relevance and value at each depth level.
  • $\theta_{3}$: [1, 2, 3, 4, 5, 6, 7, …], where the weight increases linearly with depth, emphasizing value as the trie deepens.
  • $\theta_{4}$: [1, 2, 4, 8, 16, 32, 64, …], representing an exponential increase in weight with depth, further prioritizing value.
  • $\theta_{5}$: [2, 4, 8, 16, 32, 64, 128, …], offering a more pronounced exponential tendency compared to $\theta_{4}$.
These configurations allow for nuanced control over the relevance-value trade-off, enabling the system to adapt its focus based on the hierarchical structure of the weighted trie. The choice of $\theta_{d}$ thus plays a pivotal role in optimizing the overall performance of the inference system.

### v2

@@
 Abstract.
-In the realm of sponsored search advertising, matching advertisements with the search intent of a user’s query is crucial.
-Query-to-bidwords(i.e.
-bidding keywords) rewriting, which involves transforming user queries into keywords for bidding, is a vital technique that has garnered significant attention from both industry and academia.
-Recently, with the prevalence of large language models (LLMs), generative retrieval methods have proven effective in producing high-relevance rewrites.
-However, we have identified a significant limitation in existing approaches: While fine-tuning LLMs for specific domains enhances semantic relevance, these models have no perception of the intrinsic value of their generated outputs, such as commercial value.
-Therefore, after supervised fine-tuning (SFT), a reinforcement learning from human feedback (RLHF) phase is often employed to address this issue.
-Nevertheless, traditional preference alignment methods often face challenges in aligning fine-grained values and are susceptible to overfitting, which diminishes the effectiveness and quality of the generated results.
-To address these challenges, we propose VALUE (Value-Aware Large language model for qUery rewriting via wEighted trie), the first framework that ensures the generation of high-value and highly relevant bidwords.
-Our approach utilizes weighted trie, an innovative modification of the traditional trie data structure.
-By modulating the LLM’s output probability distribution with value information from the trie during decoding process, we constrain the generation space and guide the trajectory of text production.
-Our method not only addresses fine-grained value alignment but also effectively reduces the hallucination issues often encountered with LLMs.
-Offline experiments demonstrate the effectiveness of our method in semantic matching and preference alignment, showing a remarkable improvement in the value attribute by more than fivefold.
+Query-to-bidword (i.e., bidding keyword) rewriting is fundamental to sponsored search, transforming noisy user queries into semantically relevant and commercially valuable keywords.
+Recent advances in large language models (LLMs) improve semantic relevance through generative retrieval frameworks, but they rarely encode the commercial value of keywords.
+As a result, rewrites are often semantically correct yet economically suboptimal, and a reinforcement learning from human feedback (RLHF) stage is usually added after supervised fine-tuning (SFT) to mitigate this deficiency.
+However, conventional preference alignment frequently overemphasize the ordering of bidword values and is susceptible to overfitting, which degrades rewrite quality.
+In addition, bidword value changes rapidly, while existing generative methods do not respond to these fluctuations.
+To address this shortcoming, we introduce VALUE (Value-Aware Large language model for qUery rewriting via wEighted trie), a framework that integrates value awareness directly into generation and enhances value alignment during training.
+VALUE employs the Weighted Trie, a novel variant of the classical trie that stores real-time value signals for each token.
+During decoding, the framework adjusts the LLM’s token probabilities with these signals, constraining the search space and steering generation toward high-value rewrites.
+The alignment stage uses a fine-grained preference learning strategy that emphasizes stable, high-value differences and down-weights noisy or transient fluctuations, thereby improving robustness and reducing overfitting.
+Offline experiments show that VALUE significantly outperforms baselines in both semantic matching and value-centric metrics.
 Online A/B tests further revealed that our Revenue Per Mille (RPM) metric increased by 1.64%.
@@
 Query Rewrite; Semantic Matching; Generative Retrieval; Large Language Model;
-These authors contributed equally to this work and should be considered co-first authors.
-This work was done when the first author was an intern at Alimama.
+This author contributed to this work during his internship at Alimama.
 copyright: acmlicensedjournalyear: 2024conference: Make sure to enter the correct conference title from your rights confirmation emai; TBD; TBDisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Query reformulationccs: Information systems Top-k retrieval in databasesccs: Computing methodologies Natural language processing
 1. Introduction
+Sponsored search is a crucial element of modern search engines (Fain and Pedersen, 2006).
+Its effectiveness depends on accurately mapping user queries to relevant, high-value bidwords auctioned by advertisers.
+Query texts, however, are linguistically heterogeneous: they may contain imprecise phrasing, spelling errors, or synonyms.
+These variations obscure user intent and reduce matching accuracy.
+For instance, the query “drinks for fat person” should be reformulated as “low-sugar healthy drinks” before it can align with advertiser keywords.
+Meanwhile, the rapid expansion of e-commerce has further diversified query patterns.
+Current search behavior includes long-tail expressions, ambiguous product descriptions, and multi-intent requests, each linked to distinct advertisement sets with substantially different commercial value.
+Addressing these challenges requires query rewriting and keyword matching frameworks that incorporate both semantic equivalence and value awareness.
+Such systems must improve intent recognition while simultaneously optimizing for the expected value of candidate bidwords.
+In recent years, generative query rewriting has garnered significant attention, with most state-of-the-art approaches following a three stage paradigm comprising supervised fine-tuning (SFT), value alignment, and Trie-constrained decoding (Peng et al., 2023; Chen et al., 2025; Wang et al., 2024a; Liu et al., 2025).
+In the SFT stage, domain-specific knowledge is injected into a large language model (LLM) through carefully curated rewriting corpora.
+The model then leverages its powerful semantic understanding, reasoning, and summarization capabilities to transform complex, noisy user queries into semantically canonical forms.
+During value alignment, reinforcement learning methods such as Proximal Policy Optimization (PPO), Direct Preference Optimization (DPO), Knowledge Transfer Optimization (KTO), and ORPO (Ouyang et al., 2022; Rafailov et al., 2024; Ethayarajh et al., 2024; Hong et al., 2024) adjust the model’s preferences so that the generated rewrites maximize not only semantic fidelity but also economic objectives including click-through rate and conversion likelihood.
+In the final generation stage, Trie-based decoding limits the output space to valid keyword candidates, ensuring syntactic correctness and retrieval feasibility.
+Historically, research has largely concentrated on better knowledge injection during SFT and on incorporating relevance, fluency, and factual consistency during alignment.
+Despite these advances, current systems remain inadequate in value awareness:
+  1. (1)
+    A fundamental misalignment exists between the commercial value ranking of keywords (driven by bid price and conversion potential) and their semantic relevance ranking in query rewriting.
+This divergence implies that semantically optimal rewrites rarely coincide with maximal-value keywords, creating inherent tension for generative rewriting systems.
+Recent frameworks like MoBGM (Liu et al., 2025) address this by jointly optimizing value and fidelity during alignment, yet require delicate balancing of training dynamics (e.g., reward weighting) to prevent objective dominance, which significantly increasing complexity.
+  2. (2)
+    Dynamic bid adjustments further cause real-time keyword value fluctuations.
+Consequently, value rankings observed during training become obsolete at inference, creating a temporal mismatch that undermines models relying on static/ historical signals.
+Current systems thus lack real-time value awareness, limiting adaptability to market dynamics and compromising economic optimality.
 Figure 1.
 System Overview of our proposed VALUE Framework includes three parts: ”SFT and Alignment”, ”Weighted Trie Construction and Weight Momentum Update”, and ”Collaborative Inference”.
-During the LLM decoding process, we derive a value distribution from a weighted trie to constrain the generation space, and subsequently merge this with another distribution to produce the final output distribution.
-[…]
-Sponsored search is a crucial element of modern search engines, where advertisers bid on relevant bidwords to display their ads alongside original search results.
-Upon receiving user’s query, it is essential to match the query with relevant bidwords (Fain and Pedersen, 2006), and subsequently identify suitable ads associated with those bidwords.
-This process is a critical component in the ad retrieval pipeline, driving the major revenue of search engines (Jansen and Mullen, 2008).
-The success of sponsored search depends on the search engine’s ability to accurately identify bidwords that closely align with user intent.
-However, users often face challenges when formulating queries, leading to suboptimal search experiences.
-These challenges are intensified when users do not use precise or correct terminology, employ synonyms, or mix languages in their search phrases due to varying levels of language proficiency.
-Additionally, search terms might be misspelled or overly general, complicating the retrieval of relevant ads.
-For example, a user might search for ”spring fashion,” which is broad and could correspond to bidwords like ”2024 spring collection” or ”women’s spring trends”.
-Each query reflects different intents but lacks clarity without additional context.
-As e-commerce platforms continue to grow in both scale and diversity, ensuring accurate and relevant ad retrieval becomes increasingly challenging, necessitating the need for advanced query rewriting and bidword matching techniques to better align user queries with advertiser bidwords.
-In query rewriting, two primary paradigms exist: discriminative methods, which reformulate queries by the retrieval of similar terms and generative methods, which utilize language models for direct query-to-bidword transformation.
-Discriminative methods leverage sparse and dense retrieval techniques to find relevant bidwords.
-Sparse retrieval techniques, such as BM25 (Robertson et al., 2009), employ exact lexical matching through high-dimensional term-frequency vectors.
-Their reliance on exact lexical matching leads to semantic brittleness.
-This rigidity prevents them from recognizing semantic equivalences between different expressions, resulting in significant retrieval failures when faced with semantically similar but lexically divergent queries.
-Dense retrieval approaches overcome the limitations of sparse methods by utilizing low-dimensional semantic embeddings generated by neural encoders like sBERT (Reimers, 2019).
-Dense retrieval, while effective in capturing semantic relationships, suffers from significant limitations, particularly the Matthew Effect and long-tail challenges.
-High-frequency queries dominate model training due to abundant click logs, leading to over-optimized embeddings for head queries while neglecting tail queries with sparse data.
-Furthermore, dense retrieval models struggle with distribution shifts and require costly strategies like meta-learning or data augmentation to address long-tail issues (Peng et al., 2023).
-Recently, the rise of generative retrieval has garnered widespread attention.
-Wang (Wang et al., 2024) utilizes LLMs for one-step query-to-keyword rewriting, using a trie during the decoding process.
-Despite improvements over traditional information retrieval (IR) baselines, existing generative approaches using LLMs for query-to-bidword rewriting tasks have notable limitations.
-A key limitation is their inability to recognize the value of the generated results, which, in our scenario, pertains to the outcomes they can deliver.
-Models trained solely through supervised fine-tuning on online logs lack awareness of the value of the results.
-Due to the training methodology and the SFT loss function design, the model treats each query-bidword pair in the training samples equally, failing to learn any preference or value information.
-Although there is a great deal of work on Reinforcement Learning from Human Feedback (RLHF), such as PPO, DPO, KTO, and ORPO(Ouyang et al., 2022; Rafailov et al., 2024; Ethayarajh et al., 2024; Hong et al., 2024), which can partially mitigate the value perception issues of SFT models, these methods also have significant drawbacks.
-Firstly, they require substantial computational resources for training.
-However, the effective Cost Per Mille (eCPM) value of bidwords is a statistical measure that frequently fluctuates.
-When these fluctuations occur, the previous alignment efforts become obsolete, and the model needs to be retrained.
-Secondly, these training methods are prone to overfitting (Meng et al., 2024), especially in our scenario.
-Unlike traditional alignment objectives for LLMs, such as harmlessness and usefulness, which can be achieved using a small dataset, the value of a bidword is completely independent of its textual content.
-Therefore, the model must be trained individually for each bidword.
-Our experiment on DPO show that models trained in this manner tend to overfitting.
-There is an urgent need for a new paradigm to address the challenges above.
-Through our observations, we have identified an opportunity in the sponsored search advertising scenario to enhance the value of bidwords without compromising their relevance.
-The value of a bidword does not necessarily correlate with its literal meaning.
-Two bidwords might differ by only a symbol or character, yet exhibit substantial differences in value.
-Therefore, there is an opportunity to increase the value of the top $K$ bidwords while minimally affecting relevance by prioritizing those with the same level of relevance but higher value over others.
-Thereby, we present VALUE, an innovative reward-guided inference framework that integrates value-awareness into the decoding process via a novel weighted trie structure.
-We design algorithms for the construction and update of the value of trie node.
-During decoding process, VALUE obtain reward information from for the next tokens the weighted trie.
-This information is used to adjust the output distribution, thereby sampling the subsequent token.
-By this reward-guided search, the framework aims to maximize the reward of the outputs.
-We used a multi-task fine-tuning strategy to equip the model with essential query rewriting skills.
-To achieve better value alignment, we designed an optional post-training alignment phase.
-To evaluate the effectiveness of our proposed method, we conducted offline experiments using historical online logs.
-Compared to the current online service model, our method demonstrated a 9.3% improvement in relevance.
-Additionally, online A/B tests showed increases in key profit metrics, with Cost rising by 1.53% and Revenue Per Mille (RPM) increasing by 1.64%.
-Given these significant improvements, our method has been fully deployed in the online service.
+During the LLM decoding process, we derive a value distribution from a Weighted Trie to constrain the generation space, and subsequently merge this with another distribution to produce the final output distribution.:
+To address the aforementioned challenges, we propose shifting the focus of value integration from the training phase to the generation phase, where bidword value can be dynamically incorporated without sacrificing semantic relevance.
+We introduce a novel reward-guided inference framework, named as VALUE, and design a lightweight Weighted Trie structure that embeds bidword value directly into the decoding process, as illustrated in Figure 1.
+Specifically, during Trie construction, we annotate each node with downstream bidword value information, providing token-level commercial signals along the generation path.
+This allows the model to reason over both linguistic plausibility and economic return at each decoding step.
+To account for real-time bid fluctuations, we further develop a dynamic Trie update strategy that refreshes value annotations on the fly, ensuring the system remains responsive to changes in keyword valuation during inference.
+At each decoding step, the model queries the Weighted Trie for value-based rewards and modulates the token probability distribution accordingly.
+By maximizing cumulative output value during inference, this process decouples economic optimization from semantic learning, mitigating catastrophic forgetting and overfitting associated with value-centric training objectives.
+The framework therefore supports an explicit trade off between rewriting relevance and commercial utility.
+To further bias generation toward consistently high-value bidwords, we refine DPO with adaptive loss weighting.
+The revised loss amplifies preference pairs exhibiting large, stable value gaps while down-weighting pairs subject to transient noise.
+This selectivity reduces sensitivity to unstable signals and improves generalization to long-tail and dynamically priced keywords.
+Overall, VALUE achieves a balanced integration of semantic fidelity and economic efficiency, bridging the gap between intent understanding and revenue maximization in generative query rewriting.
+To evaluate the effectiveness of the proposed method, we conduct comprehensive offline experiments on historical online logs.
+Compared with the currently deployed production system, our approach achieves a 9.3% improvement in semantic relevance.
+Furthermore, large-scale online A/B testing demonstrates consistent gains in key business metrics: consumption increases by 1.53%, and revenue per thousand impressions (RPM) rises by 1.64%.
+Given these significant and sustained improvements, our method has been fully integrated into the live serving pipeline.
 Our contributions can be summarized as follows:
   1. (1)
-    Pioneering Query Rewriting Framework via Collaborative Inference: We present a groundbreaking query rewriting framework that integrates LLMs with weighted trie, uniquely designed to simultaneously optimize for both high value and semantic relevance.
-This framework represents the first of its kind to effectively balance these two critical aspects.
+    A pioneering reward-guided collaborative inference framework for query rewriting: We propose a novel generative rewriting framework that synergistically combines large language models (LLMs) with a value-augmented Weighted Trie, enabling dynamic, lightweight value awareness while preserving semantic fidelity, achieving a principled balance between relevance and commercial utility.
   2. (2)
-    Renovation of Trie: We have reinvented the trie data structure, by seamlessly integrating value information directly into its nodes.
-This advancement is further supported by our proposed algorithms for trie construction and momentum update of node value.
+    Innovative Trie structure with embedded value propagation: We redesign the classical Trie data structure by annotating each node with downstream bidword value information, providing token level commercial-value signals during generation.
+To our knowledge, this is the first work to transform the traditionally Trie-based constraint decoding into an active value-awareness mechanism within generative rewriting, allowing real-time adaptation to fluctuating bidword value through a dynamic update strategy.
   3. (3)
-    Post-Training Alignment for Fine-Grained Reward Optimization: We propose an post-training alignment algorithm that fine-tunes the model to better align with fine-grained reward attributes.
+    Fine-grained reward optimization in post-training alignment: We introduce an enhanced preference learning algorithm that selectively emphasizes bidword pairs with large and stable value gaps, while down-weighting those susceptible to transient bid noise.
+This selective focus reduces noise exposure and mitigates overfitting risks, leading to more robust and generalizable value modeling.
 2. Related Works
@@
 The process of query rewriting, often termed query expansion or reformulation, is essential in enhancing e-commerce search technologies.
-It profoundly affects the shopping experience of users and the financial performance of e-commerce platforms.
 This approach can be generally classified into two classes: discriminative methods and generative methods.
-Discriminative methods approach query rewriting as a retrieval task.
-Pseudo-relevance feedback methods (Xu and Croft, 2017b; Cao et al., 2008; Tao and Zhai, 2006; Xu and Croft, 2017a) enhance queries by identifying expansion terms from the top-ranked documents of an initial query, combining global corpus analysis with local feedback.
-Although these methods effectively address word mismatches, they are susceptible to semantic drift due to noisy or irrelevant top results.
-To address these challenges, researchers suggest utilizing a well-constructed thesaurus as a candidate set for query rewriting.
-(Bhogal et al., 2007; Mandal et al., 2019) However, they caution that the effectiveness of these methods is highly dependent on the thesaurus’s quality.
-If the thesaurus is inadequate, it may lead to query semantic drift, where the intended meaning of the query is compromised.
-Additionally, some approaches (Antonellis et al., 2008; Cui et al., 2002; Li et al., 2022) focus on generating candidate rewrites by leveraging search logs, using similar terms from users’ search histories as extensions.
-However, these search logs inherently display a bias towards popular queries due to the Matthew effect.
-Consequently, the training data derived from this method may fall short in optimizing for less frequently searched long-tail queries.
-Generative methods, on the other hand, focus on the direct transformation of queries into keywords (Wang et al., 2024; Lee et al., 2018; Qi et al., 2020; Mohankumar et al., 2023).
-These methods involve transforming the input query directly into multiple bidwords, utilizing auto-regressive language models combined with trie-based decoding techniques to restrict the output space during inference.
-However, while previous methods ensure semantic relevance between input and output, they fall short in addressing value—a critical attribute in sponsored search advertising—which remains an open challenge.
+Discriminative methods frame query rewriting as a retrieval task.
+Pseudo-relevance feedback techniques (Xu and Croft, 2017a; Cao et al., 2008; Tao and Zhai, 2006; Xu and Croft, 2017b) expand queries by extracting terms from top-ranked initial documents, blending global corpus statistics with local feedback.
+While effective for lexical mismatch, these methods remain vulnerable to semantic drift from noisy top results.
+To mitigate this, curated thesauri have been proposed as candidate sources (Bhogal et al., 2007; Mandal et al., 2019), though their efficacy critically depends on thesaurus quality.
+Alternate approaches (Antonellis et al., 2008; Cui et al., 2002; Li et al., 2022) leverage search logs to generate rewrites, using historically similar terms as expansions.
+However, frequency bias in click data skews training toward head queries, yielding over-optimized embeddings that neglect long-tail queries.
+Dense retrieval models further struggle with distribution shifts, requiring costly mitigation strategies (e.g., meta-learning (Peng et al., 2023)) for long-tail generalization.
+Consequently, such methods inherently underperform for infrequent queries.
+Generative methods, on the other hand, focus on the direct transformation of queries into bidwords (Wang et al., 2024b; Lee et al., 2018; Qi et al., 2020; Mohankumar et al., 2023; Wang et al., 2024a; Liu et al., 2025; Chen et al., 2025).
+Most of them adopt a standard pipeline of supervised fine-tuning, value alignment, and Trie-based decoding, which constrains outputs to bidwords that are both semantically relevant and commercially valuable.
+Earlier researches have improved supervised fine-tuning through prompt engineering to better exploit LLM reasoning and summarization, and has extended value alignment to incorporate objectives such as relevance and factual consistency.
+Nonetheless, existing systems still struggle to balance semantic fidelity with economic value and fail to adapt to real-time bidword fluctuations.
 Our approach, as detailed in the following sections, effectively tackles this challenge by introducing innovative techniques that not only ensure semantic relevance but also optimize for value.
-By addressing this critical gap, our research has the potential to revolutionize sponsored search advertising, leading to more effective ad targeting and increased user satisfaction.
 2.2. Preference Alignment
-Reinforcement Learning from Human Feedback (RLHF) represents a pivotal technique for aligning LLMs with human preferences and values (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022), commonly applying the Bradley-Terry model (Bradley and Terry, 1952) to estimate the probability of a pairwise competition between two independently evaluated instances.
-The conventional RLHF pipeline is structured into three distinct phases: supervised fine-tuning (Zhou et al., 2024; Taori et al., 2023; Conover et al., 2023; Ding et al., 2023), reward model training (Gao et al., 2023; Chen et al., 2024; Havrilla et al., 2024b; Leike et al., 2018), and policy optimization (Schulman et al., 2017; Anthony et al., 2017).
-Notably, Proximal Policy Optimization (PPO) (Schulman et al., 2017) is extensively employed during the policy optimization phase.
-The RLHF framework has found broad applicability across various domains, including the mitigation of toxicity (Amini et al., 2024; Korbak et al., 2023), the assurance of safety (Dai et al., 2023), the enhancement of helpfulness (Wei et al., 2024) and the improvement of model reasoning capabilities (Havrilla et al., 2024a).
-Recent studies (Casper et al., 2023) have underscored the challenges inherent in the entire RLHF pipeline, spanning from the collection of preference data to the training of models.
-Furthermore, subsequent research has illuminated potential biases introduced by RLHF, such as the generation of verbose outputs (Dubois et al., 2024; Singhal et al., 2023; Wang et al., 2023).
+Reinforcement Learning from Human Feedback (RLHF) aligns LLMs with human preferences by applying the Bradley–Terry model (Bradley and Terry, 1952) to estimate pairwise preferences (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022).
+A typical RLHF pipeline comprises three stages: supervised fine tuning (Zhou et al., 2024; Taori et al., 2023; Conover et al., 2023; Ding et al., 2023), reward model training (Gao et al., 2023; Chen et al., 2024; Havrilla et al., 2024b; Leike et al., 2018), and policy optimization (Schulman et al., 2017; Anthony et al., 2017).
+Proximal Policy Optimization (PPO) (Schulman et al., 2017) is widely used in the final stage.
+RLHF has been applied to reduce toxicity (Amini et al., 2024; Korbak et al., 2023), improve safety (Dai et al., 2023), enhance helpfulness (Wei et al., 2024), and strengthen reasoning (Havrilla et al., 2024a).
+However, challenges arise across the pipeline, from preference data collection to model training (Casper et al., 2023), and RLHF can introduce biases such as excessively long outputs (Dubois et al., 2024; Singhal et al., 2023; Wang et al., 2023).
+Alternative alignment methods avoid explicit reward model training.
+Direct Preference Optimization (DPO) (Rafailov et al., 2024), Kahneman Tversky Optimization (KTO) (Ethayarajh et al., 2024), Simple Preference Optimization (SimPO) (Meng et al., 2024), and Monolithic Preference Optimization (ORPO) (Hong et al., 2024) optimize directly on preference pairs.
+SimPO and ORPO further remove the need for a reference model, simplifying training.
+Although these methods cannot sample preference pairs from an optimal policy, researchers augment data with outputs from a supervised fine-tuned model and employ iterative frameworks that periodically update the reference.
+Alignment as Reward Guided Search (ARGS) (Khanov et al., 2024) integrates alignment into decoding instead of training.
+Despite their advantages, these alignment techniques remain computationally intensive and must be retrained whenever value criteria change.
+RLHF methods are commonly employed in generative query rewriting to infuse bidwords value information.
+However, a weak correlation between bidwords value ranking and relevance ranking often arises during rewriting.
+This misalignment leads to overfitting and degraded rewriting relevance in value-aware generative models.
+Our proposed VALUE framework addresses these issues through a dual-pronged approach: (1) leveraging the Weighted Trie for lightweight, real-time value awareness during generation, and (2) mitigating catastrophic forgetting of SFT-injected knowledge via optimized value preference alignment.
 3. Methodology
-Traditional methods, such as standard beam search or top-k sampling, prioritize candidates with higher probabilities, often neglecting their value attributes.
-Although these approaches are effective in numerous contexts, they may fall short in applications where the value attributes are important.
-VALUE addresses this limitation by incorporating a Weighted Trie that allows for value-aware token selection.
-At each step of the decoding process, VALUE modifies the model’s output probabilities based on the value obtained from the trie, prioritizing tokens of higher value among those already highly relevant.
 Figure 2.
@@
 Then, we sample the next token from the modified distribution.
-3.1. System Overview
-As can be seen from Figure 1, our entire system comprises the following key modules: ”LLM SFT and Alignment,” ”Weighted Trie Construction and Weight Momentum Update,” and ”Collaborative Inference.” For each generated token, we obtain the probability distribution from the LLM.
-Concurrently, we retrieve all possible candidates under the current prefix from the weighted trie along with their respective values.
-By integrating these values with the probability distribution, we derive a new probability distribution and sample the next token from it.
+As illustrated in Figure 1, VALUE operates in two primary phases: (1) a training phase integrating Supervised Fine-Tuning (SFT) with value preference alignment, and (2) a collaborative inference phase utilizing the Weighted Trie structure.
+During the SFT phase, preprocessed and relevant query-bidword pairs are employed to inject domain-specific rewriting knowledge for sponsored search into LLMs.
+Subsequently, we present our lightweight value-aware inference framework and the optimized value preference alignment method in detail.
+We first introduce the construction of Weighted Trie and weight momentum update.
+Then, we will introduce collaborative inference based on Weighted Trie and the post-training alignment method we proposed in training phase.
+Finally, we will describe how VALUE is deployed for online serving.
 3.1. Weighted Trie
 3.1.1. Definition and Construction of Weighted Trie
-Our Weighted Trie is an renovation of the traditional trie data structure.
+Weighted Trie is an renovation of the traditional trie data structure.
 Each node in the Weighted Trie, referred to as a WeightedTrieNode, contains additional attributes to store weight-related information.
 The primary attributes of a WeightedTrieNode are:
-  • children: A dictionary where the keys are integers representing tokens, and the values are child nodes of type WeightedTrieNode.
-  • mean: A floating-point number representing the average weight of the node’s children.
-  • max: A floating-point number representing the maximum weight among the node’s children.
+  • children: A dictionary where the keys are integers representing tokens in LLM’s vocabulary, and the values are child nodes of type WeightedTrieNode.
+  • mean_v: A floating-point number representing the average eCPM value of the node’s children, where eCPM (effective Cost Per Mille) is the bidword value calculated offline.
+  • max_v: A floating-point number representing the maximum eCPM value among the node’s children.
   • is_word: A boolean value indicating whether the node represents a complete word.
-Construction of Weighted Trie: the construction of the Weighted Trie involves initializing the root node, inserting bidwords with their corresponding values, and updating the weights of the nodes.
+The construction of the Weighted Trie involves initializing the root node, inserting bidwords with their corresponding values, and updating the weights of the nodes.
 The pseudocode in Algorithm 1 demonstrates the construction process.
 Algorithm 1 Construct Weighted Trie
-[…]
+Input: List of pairs $(\textit{bidword},\textit{ecpm})$
 Output: Weighted Trie
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+ Initialize weighted trie: $\mathit{weighted\_trie}$
+for each $(\textit{bidword},\textit{ecpm})$ in list do
+  $\mathit{token\_ids}\leftarrow\text{tokenizer}(\textit{bidword})$
+  Insert $\mathit{token\_ids}$ into the trie:
+  for each $\mathit{token\_id}$ in $\mathit{token\_ids}$ do
+   Traverse or create node for $\mathit{token\_id}$
   end for
-[…]
-[…]
+  Set $\mathit{leaf\_node.mean\_v}\leftarrow\textit{ecpm}$
+  Set $\mathit{leaf\_node.max\_v}\leftarrow\textit{ecpm}$
 end for
@@
 for each node in post-order do
-[…]
-[…]
+  $\mathit{node.mean\_v}\leftarrow\frac{1}{N}\sum_{\mathit{child}\in\mathit{children}}\mathit{child.mean\_v}$
+  $\mathit{node.max\_v}\leftarrow\max_{\mathit{child}\in\mathit{children}}\mathit{child.max\_v}$
 end for
 3.1.2. Weighted Trie Momentum Update
-Momentum updates are essential due to the frequent changes in eCPM and the evolving bidword space. The process is delineated as follows:
+Momentum updates are essential due to the frequent changes in bidding prices and the evolving bidword space. The process is delineated as follows:
 Tokenization and Node Insertion:
 Each bidword undergoes tokenization, converting it into a sequence of token IDs. These IDs facilitate the insertion or updating of nodes within the trie. The update process for a node is governed by:
-(1) $V_{\text{new}}=\max(eCPM_{\text{new}},V_{\text{old}})$
-(2) $W_{\text{new}}=\alpha\cdot eCPM_{\text{new}}\ +\ \beta\cdot W_{\text{old}}$
-where $\alpha$ and $\beta$ are hyperparameters that control the update rate, with the constraint $\alpha+\beta=1$.
-V is the max attribute of the node and W is the mean attribute of the node.
-If the node is absent, it is initialized as:
-(3) $W_{\text{new}}=eCPM_{\text{new}},\ V_{\text{new}}=eCPM_{\text{new}}$
+(1) $Max\_V_{\text{new}}=\max(eCPM_{\text{new}},Max\_V_{\text{old}})$
+(2) $Mean\_V_{\text{new}}=\alpha\cdot eCPM_{\text{new}}\ +\ (1-\alpha)\cdot Mean\_V_{\text{old}}$
+where $\alpha$ is the hyperparameter that control the update rate. If the node is absent, it is initialized as:
+(3) $Max\_V_{\text{new}}=eCPM_{\text{new}},\ Mean\_V_{\text{new}}=eCPM_{\text{new}}$
 Bottom-up Update:
@@
 3.2. Collaborative Inference
-As shown in Figure 2, when generating the $t$-th token, we insert the query into the prompt template and concatenate it with the bidword prefix to form the input of LLM to get the output probabilities.
+As shown in Figure 2, when generating the $t$-th token, we insert the query into the prompt template and concatenate it with the bidword prefix(i.e.
+the first t-1 tokens) to form the input of LLM to get the output probabilities.
 We use generated bidword prefix to retrieve all child nodes from the trie.
 The value of these child nodes is computed using the following formula:
-(4) $\mathcal{V}_{k}=\alpha\cdot\text{mean}_{k}+\beta\cdot\text{max}_{k}$
-Here, $\alpha$ and $\beta$ are tunable hyperparameters, both set to 0.5, allowing us to prioritize paths with higher average values or to explore bidwords with the maximum potential value, depending on our specific needs.
+(4) $\mathcal{V}_{k}=\beta\cdot Max\_V_{k}+(1-\beta)\cdot Mean\_V_{k}$
+where $\beta$ is the tunable hyperparameter set to 0.5, allowing us to prioritize paths with higher average values or to explore bidwords with the maximum potential value, depending on our specific needs.
 We then apply the softmax function to these values to obtain the normalized value $\hat{\mathcal{V}}_{k}$.
-Next, we mask out all probabilities in the LLM-generated distribution that are not children of the current node.
+Next, we mask out all probabilities in the LLM-generated distribution that are not children of the previous generated token in the Weighted Trie.
 For the remaining probabilities, we adjust them by weighting with the normalized value $\hat{\mathcal{V}}_{k}$ and a depth-dependent factor $\theta$.
 The adjusted probability for each token $k$ is given by:
-(5) $p(k\mid x_{<t})=p_{\text{LLM}}(k\mid x_{<t})\cdot(1+\mathcal{V}_{k}\cdot\theta)$
-where $\theta$ is designed to vary with the depth of the trie, playing a crucial role in balancing relevance and value in our model’s output.
-At shallower levels of the trie, $\theta$ is set to smaller values to minimize the influence of value on the output, thereby preventing the model from favoring high-value but less relevant options.
-As we delve deeper into the trie, $\theta$ increases, allowing us to prioritize higher-value candidates from a pool of already highly relevant choices.
-Consequently, we derive a modified probability distribution $p(k\mid x_{<t})$, from which the next token is sampled as usual.
-3.4. Multi-task Fine-Tuning
-As no open-source LLMs are trained to perform query rewriting for the e-commerce scenario, they lack an awareness of the bidword space.
-Consequently, they are unable to effectively generate bidwords.
-To address this, we need to perform supervised fine-tuning (SFT) on these models to inject domain-specific knowledge into the model and enhancing their capability to understand and effectively rewrite e-commerce queries for improved user experience and search accuracy.
-Query Rewriting Dataset: To construct our training dataset, we initially obtain query-bidword pairs from online logs.
-Specifically, when a user searches for a query $x$ on our e-commerce platform, the system logs the corresponding bidwords $Y=\{y^{1},y^{2},\ldots,y^{n}\}$ that are associated with the advertisements the user interacts with.
-From this list, we select the top-ranked bidword $y^{1}$ as the gold standard candidate to construct our initial dataset $D$ with $N$ samples:
-(6) $D=\{(x_{i},y_{i}^{1})\}_{i=1}^{N}\quad\text{where}\quad x_{i}\sim p(x)$
-where $p(x)$ denotes the query distribution in our search engine.
-We apply a relevance filter to $D$ using a representation model fine-tuned from an open-source embedding model.
-The relevance score $\text{rel}(x,y)$ is computed as the cosine similarity between the embeddings of $x$ and $y$:
-(7) $\text{rel}(x,y)=\frac{\text{emb}(x)\cdot\text{emb}(y)}{\|\text{emb}(x)\|\|\text{emb}(y)\|}$
-We then filter the dataset $D$ to obtain $D_{r}$ ensuring that the relevance score exceeds a threshold $\tau_{\text{rel}}$:
-(8) $D_{r}=\{(x_{i},y_{i})\mid(x_{i},y_{i})\in D,\,\text{rel}(x_{i},y_{i})>\tau_{\text{rel}}\}$
-For head queries, the number of corresponding bidwords can be in the tens of thousands. To ensure class balance, we truncate the bidwords for each query based on their value.
-Fine-tuning Tasks: We employ two types of supervised fine-tuning (SFT) tasks to enhance the model’s understanding and performance in query rewriting for e-commerce.
-Task 1: <prompt + query + bidword>
-This task aims to enhance the model’s spatial awareness of the bidword space and the association between queries and bidwords. Subsequently, the model will develop fundamental rewriting skills.
-Task 2: <prompt + query + bidword_list>
-In this task, the model is trained with lists of bidwords sorted in descending order based on their eCPM values. The bidword_list for each query is structured as follows:
-[…]
+(5) $p_{WT}(k\mid x_{<t})=p_{{LLM}}(k\mid x_{<t})\cdot(1+\hat{\mathcal{V}}_{k}\cdot\theta_{t})$
+The fusion value $p_{WT}(k\mid x_{<t})$ is an affine transformation of $p_{{LLM}}(k\mid x_{<t})$ and $\hat{\mathcal{V}}_{k}$, selected through a series of tests and comparisons with additive fusion, multiplicative fusion, etc.
+where $\theta_{t}\in\Theta$ is designed to vary with the depth $t$ of the trie, playing a crucial role in balancing relevance and value in our model’s output.
+Specifically, $\Theta$ assigns lower value weight to initial generated tokens and higher weight to later generated tokens.
+This design stems from two key observations: (1) Initial tokens primarily establish semantic relevance.
+Excessive value focus at this stage risks compromising query intent alignment.
+(2) Subsequent tokens refine bidword specificity.
+Once core semantics are anchored, increased value weighting allows precise optimization within the relevant search space.
+Thus, $\Theta$ progressively shifts focus from relevance grounding to value maximization across the generation trajectory.
+Consequently, we derive a modified probability distribution $p_{WT}(k\mid x_{<t})$, from which the next token is sampled as usual.
 3.3. LLM Post-training Alignment with Weighted Trie
-Although conventional RLHF approaches can effectively integrate with our framework, we propose a more tailored solution.
-Our Weighted DPO (WDPO) approach directly incorporates the economic value of bidwords into the alignment process, leveraging the eCPM values for bidwords.
-It is important to note that WDPO is optional as the performance improvements it offers are not as significant relative to computational cost when compared to SFT and the varying nature of bidwords.
-Therefore, WDPO can be considered when resources permit, providing an additional layer of refinement to the training process.
+Although the Weighted Trie offers a lightweight mechanism for balancing relevance and commercial value of rewriting results, we seek to further enhance the LLM’s generation probability for high-value bidwords.
+This is critical because, in scenarios where a query has numerous relevant bidwords, high-value candidates meeting the relevance threshold might be excluded from the rewrite set due to low generation probabilities, thereby limiting the effectiveness of the Weighted Trie.
+However, prevailing value alignment techniques, exemplified by DPO, often excessively prioritize bidwords value ranking.
+This over-emphasis can lead to model overfitting or catastrophic forgetting of the semantic knowledge acquired during SFT.
+To address these challenges, we propose Weighted Direct Preference Optimization (WDPO).
+WDPO directly incorporates bidwords value into DPO, steering the model’s attention toward bidword pairs exhibiting substantial value disparities.
+This approach effectively mitigates the negative impact of the value alignment phase on rewriting relevance.
 The alignment process proceeds as follows:
@@
 3.4. Online Depolyment
+We constructed a WeightedTrie using a dataset of 30 million bidwords.
+In real-world business scenarios, the growth of the bidwords is typically gradual and has negligible impact on query processing performance.
+As an alternative, a pruning strategy based on key metrics such as eCPM or historical number of clicks can be employed to maintain the bidword count within predefined performance bounds.
 In the context of deploying LLMs for online serving, the balance between computational efficiency and latency is paramount.
-Our approach leverages the Qwen2-7B model for offline inference, targeting head queries that constitute 50% of the page views (PV) in our system.
-This offline process generates the top 500 results per query, which are subsequently cached for rapid access.
+Our approach leverages the Qwen2-7B(Team, 2024) model for offline inference, targeting head queries that constitute 85% of the page views (PV) in our system.
+This offline process generates the top 500 bidwords per query, which are subsequently cached for rapid access.
 For mid-tail and long-tail queries that do not benefit from the cached results, we employ the Qwen2-1.5B model in an online serving capacity.
-This model is optimized to deliver a latency of 50 milliseconds, efficiently producing the top 50 results per query.
+This model is optimized to deliver a latency of 50 milliseconds, efficiently producing the top 50 bidwords per query.
 In the future, we plan to integrate user behavior sequence data into our online serving framework to enhance personalized recommendation systems and gradually remove the offline part.
@@
 4.1. Datasets
-[…]
+Training Dataset: For multi-task fine-tuning, we extracted 150 billion records from 30 days of online logs, followed by several rounds of data cleaning.
+Initially, we employed manual observation and regular expressions to filter out the majority of noisy queries.
+In the subsequent screening phase, we calculated the click-through rate (CTR) based on page views (PV) and clicks, retaining only those $\langle\texttt{query},\texttt{bidword}\rangle$ pairs associated with purchase behavior that meet the click-through rate requirements.
+We then applied a relevance model for further filtering.
+This relevance model is a fine-tuned version of BAAI/bge-large-zh-v1.5(Xiao et al., 2024), optimized for e-commerce scenarios.
+We truncated the number of bidwords for each query, retaining a maximum of 50 bidwords per query.
+After third round of filtering, we were left with 110 million records.
 Test Dataset: Our test dataset was constructed by randomly sampling 30,000 queries from online logs.
@@
     Hit Rate: A precision-oriented metric measuring the alignment between generated rewrites and user-clicked bidwords. For a query set $\mathcal{Q}$, it is defined as:
-[…]
+    (10) $\text{hitrate@}K=\frac{\sum_{q\in\mathcal{Q}}|\mathcal{R}_{q}^{K}\cap\mathcal{C}_{q}|}{\sum_{q\in\mathcal{Q}}|\mathcal{C}_{q}|}$
     where $\mathcal{R}_{q}^{K}$ represents the top-$K$ rewrite candidates for query $q$, and $\mathcal{C}_{q}$ denotes the set of clicked bidwords for $q$.
@@
 4.3. Implementation Details
-Multi-task Fine-tuning: We fine-tuned the Qwen2-7B model using a learning rate of 1e-5 and a cosine learning rate scheduler.
+Supervised Fine-tuning: We fine-tuned the Qwen2-7B model using a learning rate of 1e-5 and a cosine learning rate scheduler.
 The process was optimized with AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and a weight decay of 0.001.
@@
 The training was executed on 16 NVIDIA H20 GPUs for a duration of 3 hours.
+Construction and Update of Weighted Trie: We built a Weighted Trie containing 30 million bidwords with an average length of 7.9 characters, which occupies approximately 2.7 GB of memory after serialized.
+Single-threaded construction finishes in 56 minutes on a 32-core CPU machine, but multi-threaded optimization reduces this time to roughly 20 minutes.
+We used a message queue for streaming updates to the Weighted Trie, achieved minute-level update latency.
 4.4. Offline Experiments
@@
 The eCPM metric is used as an approximate estimation of reward, serving as a reference for the effectiveness of Weighted Trie collaborative inference.
-Model hitrate@50 hitrate@500 Spearman’s $\rho$ Relevance OOVR eCPM
-OnlineModel 59.93% 85.41% 0.02 68.15% 0% 8676
-DeepNB 44.16% 79.82% 0.13 46.26% 0% 12852
-BEQUE 39.26% 62.85% 0.08 74.37% 38.41% 7582
-raw Qwen2-7B 12.88% 18.37% -0.03 59.81% 68.63% 2937
-SFT Qwen2-7B 38.18% 63.71% 0.02 76.33% 34.85% 7439
-SFT+Trie 59.11% 86.38% -0.01 75.31% 0% 7270
-SFT+Trie+DPO 38.61% 62.19% 0.05 69.48% 0% 28391
-SFT+WeightedTrie(VALUE) 60.37% 91.52% 0.46 74.55% 0% 59803
-SFT+WeightedTrie+DPO 50.35% 77.73% 0.23 70.18% 0% 38301
-SFT+WeightedTrie+WDPO(VALUE+WDPO) 62.67% 91.91% 0.56 73.49% 0% 63775
+Experiment Number Model hitrate@50 hitrate@500 Spearman’s $\rho$ Relevance OOVR eCPM
+Exp.1 OnlineModel 59.93% 85.41% 0.02 68.15% 0% 8676
+Exp.2 DeepNB 44.16% 79.82% 0.13 46.26% 0% 12852
+Exp.3 BEQUE 39.26% 62.85% 0.08 74.37% 38.41% 7582
+Exp.4 BEQUE+Trie 42.77% 65.36% 0.04 72.96% 0% 7225
+Exp.5 BEQUE+WeightedTrie 53.94% 79.20% 0.26 70.73% 0% 19275
+Exp.6 raw Qwen2-7B 12.88% 18.37% -0.03 59.81% 68.63% 2937
+Exp.7 SFT Qwen2-7B 38.18% 63.71% 0.02 76.33% 34.85% 7439
+Exp.8 SFT+Trie 59.11% 86.38% -0.01 75.31% 0% 7270
+Exp.9 SFT+DPO+Trie 38.61% 62.19% 0.05 69.48% 0% 28391
+Exp.10 SFT+WeightedTrie 60.37% 91.52% 0.46 74.55% 0% 59803
+Exp.11 SFT+DPO+WeightedTrie 50.35% 77.73% 0.23 70.18% 0% 38301
+Exp.12 SFT+WDPO+WeightedTrie(VALUE) 62.67% 91.91% 0.56 73.49% 0% 63775
 In this section, we present the results of offline experiments conducted to evaluate the performance of our proposed models against several baseline models.
 The comparative performance analysis is summarized in Table 1.
-We compared our models with the following baselines: first, the Online Model, which represents the current state-of-the-art multi-channel recall system deployed online; second, the DeepNB model (Chen et al., 2020), which employs a vector-based approach focusing solely on value; and finally, the BEQUE model (Peng et al., 2023), which utilizes a LLM fine-tuned with multi-task SFT and PRO alignment, representing the latest query rewriting approach.
+We compared our models with the following baselines: first, the Online Model, which represents the current state-of-the-art multi-channel recall system deployed online; second, the DeepNB model (Chen et al., 2020), which employs a vector-based approach focusing solely on bidword value; and finally, the BEQUE model (Peng et al., 2023), which utilizes a LLM fine-tuned with multi-task SFT and PRO alignment(Song et al., 2024), representing the generative query rewriting approach.
 The performance metrics used for evaluation include hitrate, eCPM, Spearman’s $\rho$, Relevance, and OOVR(Out-Of-Vocabulary Rate.
@@
 4.4.1. Comprehensive Evaluation of the VALUE Framework
-The VALUE framework we propose is designed to optimize online RPM revenue by utilizing offline eCPM values as a reference for learning effectiveness, rather than as direct revenue indicators.
-Our subsequent online experiments have demonstrated the validity and rationality of this modeling approach.The results presented in Table 1 clearly illustrate the significant effectiveness of our VALUE framework.
-Notably, the eCPM metric shows a remarkable improvement, serving as a strong indicator of offline learning success.
-Compared to the model trained with Supervised Fine-Tuning (SFT) on the same dataset, our VALUE model achieves a 722.6% increase in eCPM.
-Furthermore, when compared to the current online service model, our framework exhibits a 589.3% enhancement in eCPM.
-Importantly, the relevance metric for our VALUE model does not decline when compared to the SFT model; in fact, it demonstrates an improvement of 6.4 percentage points over the online service model.
-This observation substantiates our earlier hypothesis that relevance and value are decoupled in this domain, allowing us to reorder the generation sequence of bidwords to enhance both metrics simultaneously without trade-offs.
-Compared to the DeepNB model, which focuses on efficiency-oriented rewriting using click data to maximize value, our VALUE framework demonstrates superior performance across multiple dimensions.
-While the DeepNB model effectively maximizes value, it suffers from poor relevance metrics.
-In contrast, our VALUE framework not only maintains high relevance but also significantly outperforms DeepNB in terms of value attributes.
-Specifically, the eCPM metric for our VALUE model is substantially higher than that of DeepNB.
-When compared to the BEQUE model, our framework achieves comparable relevance metrics but excels in generating high-value bidwords.
-This is evidenced by the eCPM metric, where VALUE achieves a value of 59803, far surpassing BEQUE’s 7582.
-Additionally, our VALUE framework demonstrates a superior hit rate, with hitrate@500 reaching 91.52%, compared to BEQUE’s 62.85%.
-This significant improvement in hit rate underscores the effectiveness of our approach in generating bidwords that are not only relevant but also highly likely to be clicked by users.
-It is noteworthy that the BEQUE method suffers from a high Out-Of-Vocabulary Rate (OOVR) of 38.41%, which adversely affects its hit rate.
-Moreover, our VALUE framework outperforms the online service model across all evaluated metrics, further validating the efficacy of our approach.
-The comprehensive improvements in eCPM, relevance, and other metrics underscore the robustness and effectiveness of our method in generating high-value and contextually relevant bidwords for e-commerce query rewriting tasks.
-Figure 3. Trade-off between RPM and PV rele through $\theta$.
-[…]
+As evidenced in Table 1, VALUE (Exp.12) demonstrates superior performance across all evaluation metrics, underscoring the robustness and effectiveness of our proposed method in generating high-value and contextually relevant bidwords for e-commerce query rewriting.
+DeepNB (Exp.2) leverages click logs for efficiency-oriented rewriting to maximize value, yet its rewrites suffer from compromised relevance to the original search queries.
+Conversely, the generative approach BEQUE (Exp.3) capitalizes on the semantic comprehension capabilities of LLMs, achieving high rewrite relevance, but exhibits poor value awareness regarding bidwords.
+In contrast to both methods, VALUE significantly enhances the eCPM and hitrate of the rewritten queries while maintaining high relevance.
+This indicates that VALUE effectively synthesizes the complex semantic understanding derived from LLMs with value awareness enabled by the Weighted Trie.
+Notably, BEQUE exhibits a remarkably high Out-of-Vocabulary Word Rate (OOVR) of 38.41%, and integrating the Trie-based inference structure into BEQUE (Exp.4) significantly enhances hitrate while ensuring output authenticity.
+Furthermore, compared to the mainstream generative method ”SFT+Trie+DPO” (Exp.9), VALUE achieves improvements across all evaluation metrics.
+This outcome suggests that our proposed reward-guided inference framework mitigates potential overfitting inherent in the value alignment stage.
+4.4.2. Ablation Study of the Weighted Trie.
+Our proposed Weighted Trie guides LLMs during the generation phase to produce bidwords that optimize both relevance and value, while adapting to real-time changes in their monetary value.
+Compared to ”SFT+Trie” (Exp.8), the ”SFT+WeightedTrie” configuration (Exp.10) achieves a 5.1 percentage point increase in hitrate@500 and substantially improves eCPM while concurrently maintaining rewrite relevance.
+Similarly, when employed in the generation phase alongside other generative rewriting methods (Exp.4 vs Exp.5, Exp.9 vs Exp.11), Weighted Trie significantly enhances key evaluation metrics associated with bidwords value.
+These results demonstrate that our Weighted Trie serves as an efficient, lightweight, and pluggable module that effectively enhances the value awareness of generative query rewriting methods.
+Notably, the incorporation of Weighted Trie significantly elevates the Spearman’s $\rho$, demonstrating that its value-relevance balancing mechanism during decoding enhances fine-grained value awareness.
+4.4.3. Ablation Study of WDPO.
+Traditional preference alignment techniques employed in BEQUE, such as DPO and Preference Ranking Optimization (PRO), are susceptible to significant overfitting during training, degrading performance in terms of both hitrate and relevance.
+The overfitting issue inherent to DPO arises from its tendency to overfit specific preference pairs in the training data, leading to compromised generalization capabilities.
+Consistent with this observation, Table 1 demonstrates marked declines in both hitrate and relevance when integrating DPO with either the ”SFT+Trie” (Exp.8 vs Exp.9) or the ”SFT+WeightedTrie” (Exp.10 vs Exp.11).
+In contrast, WDPO enables LLMs to focus learning efforts on high-value bidwords by attenuating the influence of preference pairs exhibiting minimal value differences during training.
+Consequently, the VALUE framework incorporating WDPO (Exp.12) achieves substantial gains across multiple value-centric metrics while preserving relevance.
+Figure 3. Trade-off between RPM and PV rele through $\Theta$. The specific values of RPM and PV rele are confidential, so the RPM and PV rele here are equivalently scaled.:
 Table 2. Comparative Performance Analysis of Hyperparameter.
 Hyperparameter($\Theta$) hitrate@500 Relevance
-$\theta_{1}$ 86.38% 75.31%
-$\theta_{2}$ 89.65% 75.04%
-$\theta_{3}$ 91.52% 74.55%
-$\theta_{4}$ 88.63% 69.85%
-$\theta_{5}$ 73.71% 60.53%
-4.4.2. Enhanced Performance in Fine-Grained Value Alignment
-Compared to traditional alignment methods such as DPO and Preference Ranking Optimization (PRO) used in the BEQUE model, our VALUE framework demonstrates a significant lead in the Spearman rank correlation metric.
-This indicates that conventional value alignment methods struggle with fine-grained value alignment tasks.
-Specifically, the Spearman’s $\rho$ for our VALUE model is 0.46, which is substantially higher than the 0.05 achieved by DPO model and 0.08 by BEQUE.
-Additionally, our framework seamlessly integrates with conventional model alignment methods.
-When we combined our proposed WDPO with the VALUE framework, we achieved the highest Spearman’s $\rho$ of 0.56.
-DPO suffers from significant overfitting issues, which severely impact its performance in terms of hit rate and relevance.
-In our experiments, the SFT + DPO + Trie model exhibited a hitrate@500 of only 62.19% and a relevance score of 69.48%, both of which are considerably lower than those achieved by our VALUE framework.
-DPO faces inherent challenges when dealing with positive and negative samples that are lexically very similar.
-The overfitting problem in DPO arises because the method tends to excessively optimize for the specific preference pairs seen during training, leading to a decline in generalization capability.
-DPO enhances the model’s belief in specific events.
-Therefore, when we use DPO in conjunction with our VALUE framework, the improvement in eCPM is not significant.
-Specifically, the eCPM for the VALUE + DPO model is 38301, which is only marginally higher than the 28391 achieved by the SFT + DPO + WeightedTrie model.
-Furthermore, the improvement in the Spearman rank correlation metric with DPO is not significant.
-The SFT + DPO + Trie model achieved a Spearman’s $\rho$ of only 0.05, which is marginally better than the baseline but far from satisfactory.
-This limited improvement underscores the inherent limitations of DPO in handling fine-grained value alignment tasks.
+$\Theta_{1}$: [0, 0, 0, 0, 0, 0, 0, …] 86.38% 75.31%
+$\Theta_{2}$: [1, 1, 1, 1, 1, 1, 1, …] 89.65% 75.04%
+$\Theta_{3}$: [1, 2, 3, 4, 5, 6, 7, …] 91.52% 74.55%
+$\Theta_{4}$: [1, 2, 4, 8, 16, 32, 64, …] 88.63% 69.85%
+$\Theta_{5}$: [2, 4, 8, 16, 32, 64, 128, …] 73.71% 60.53%
 4.5. Exploration of Hyperparameter Configurations
-In this section, we provide a concise analysis of the impact of the hyperparameter $\theta$ on system performance.
-As shown in Table 2, an increase in $\theta$ results in a decrease in relevance.
-Figure 3 visually represents the dynamics of this trade-off.
-Initially, optimizing for value leads to an increase in RPM; however, when relevance becomes significantly low, RPM ultimately decreases.
-This underscores the critical importance of meticulous hyperparameter tuning to ensure alignment with specific application objectives.
-For a detailed exploration of the hyperparameter $\theta$, please refer to Appendix A.
+In this section, we conduct a thorough investigation into the impact of varying the hyperparameter $\Theta$ on system performance.
+This parameter, which scales with the depth of the Weighted Trie, modulates the trade-off between relevance and commercial value at different stages of the inference process.
+Specifically, we evaluate five distinct configurations (see Table 2):
+  • $\Theta_{1}$: Assigns a value weight of 0 throughout, effectively disabling the Weighted Trie and relying solely on the original LLM inference output.
+  • $\Theta_{2}$: Assigns equal weighting to relevance and value at every depth of the Weighted Trie.
+  • $\Theta_{3}$: Increases the value weight linearly as a function of Weighted Trie depth.
+  • $\Theta_{4}$: Increases the value weight exponentially with Weighted Trie depth, significantly prioritizing value in the deeper inference stages.
+  • $\Theta_{5}$: Further increases the baseline value weights at all depths established by $\Theta_{4}$.
+Table 2 and Figure 3 present the performance metrics under different $\Theta$ settings across both online and offline evaluations.
+Key observations reveal that increasing the value weight consistently compromises relevance.
+Critically, however, excessive value weighting (as in $\Theta_{4}$ and $\Theta_{5}$) not only severely degrades relevance but also diminishes RPM and hitrate.
+Within the sponsored search context, the objective of query rewriting is to maximize the value of rewrites while satisfying a minimum relevance criterion.
+The $\Theta_{3}$ configuration addresses this by directing VALUE to prioritize highly relevant bidwords during the initial token generation stages, subsequently shifting focus to high-value bidwords later in the process.
+This strategy strikes an optimal balance between relevance and value by first ensuring the generation of a more relevant prefix.
+Our findings underscore that meticulous hyperparameter tuning is paramount for aligning model behavior with specific sponsored search objectives.
 4.6. Scaling Laws in Model Performance
@@
 4.8. Case Study
-Figure 4. Case Study of VALUE Framework
-[…]
-In this section, as illustrated in Figure 4, we examine a specific case involving the user query ”fatty’s drink.” Previously, without the use of LLMs, our system could only match the query at the keyword level, specifically ”drink.” With the integration of LLMs, we gained a deeper understanding of the user’s intent, identifying a preference for low-sugar, low-calorie beverages.
-Additionally, by employing a weighted trie, we ensured that the advertisements presented were not only closely aligned with the user’s intent but also of high value.
+Figure 4. Case Study of VALUE Framework:
+In this section, as illustrated in Figure 4, we examine a specific case involving the user query ”drink for fat person”.
+Previously, without the use of LLMs, our system could only match the query at the keyword level, specifically ”drink” .
+With the integration of LLMs, we gained a deeper understanding of the user’s intent, identifying a preference for low-sugar, low-calorie beverages.
+Additionally, by employing a Weighted Trie, we ensured that the advertisements presented were not only closely aligned with the user’s intent but also of high commercial value.
 5. Conclusion
-In this study, we introduced VALUE, an innovative framework designed to enhance query rewriting in sponsored search advertising by integrating value-awareness into the generation process.
-Our approach addresses the significant challenge of producing rewrites that are not only semantically relevant but also economically valuable—a gap that existing methods often fail to bridge.
-By embedding value information directly into the LLM’s decoding process through a novel weighted trie structure, we effectively steer generation towards high-value bidwords without substantially compromising relevance.
-Comprehensive experiments underscore the effectiveness of VALUE in enhancing the economic attributes of bidword generation.
-Offline evaluations demonstrate a significant increase in eCPM, outperforming traditional methods while preserving high relevance scores.
-The improvement in Spearman rank correlation confirms the superior ranking quality achieved through our fine-grained value alignment.
-Furthermore, online A/B testing validates the real-world efficacy of VALUE, highlighting notable improvements in key performance indicators such as cost and RPM.
-Our framework’s versatility also suggests potential applications beyond sponsored search advertising, wherever the generation space is fixed and reward signals are present.
-Appendix A Hyperparameter Details
-In this appendix, we delve into the specifics of the hyperparameter $\theta_{d}$, which is designed to vary with the depth of the trie.
-The selection of $\theta_{d}$ is critical in modulating the influence of value relative to relevance as the trie depth changes.
-The configurations explored in our research are as follows:
-  • $\theta_{1}$: [0, 0, 0, 0, 0, 0, 0, …], indicating that the weighted trie is ineffective and the original LLM reasoning result is used.
-  • $\theta_{2}$: [1, 1, 1, 1, 1, 1, 1, …], implying equal weighting of relevance and value at each depth level.
-  • $\theta_{3}$: [1, 2, 3, 4, 5, 6, 7, …], where the weight increases linearly with depth, emphasizing value as the trie deepens.
-  • $\theta_{4}$: [1, 2, 4, 8, 16, 32, 64, …], representing an exponential increase in weight with depth, further prioritizing value.
-  • $\theta_{5}$: [2, 4, 8, 16, 32, 64, 128, …], offering a more pronounced exponential tendency compared to $\theta_{4}$.
-These configurations allow for nuanced control over the relevance-value trade-off, enabling the system to adapt its focus based on the hierarchical structure of the weighted trie.
-The choice of $\theta_{d}$ thus plays a pivotal role in optimizing the overall performance of the inference system.
+This study introduces VALUE, an innovative framework that addresses the critical challenge of producing rewrites that simultaneously optimize semantic relevance and commercial value, which is a critical limitation pervasive in existing methods.
+VALUE incorporates real-time value signals into LLM decoding through a novel Weighted Trie and applies WDPO-based alignment that prioritizes bidword pairs with large value gaps during training.
+These components jointly guide generation toward high-value rewrites without compromising semantic relevance.
+Extensive experiments confirm the framework’s effectiveness.
+Offline tests yield substantial eCPM gains while preserving relevance, and online A/B experiments report significant improvements in cost and Revenue Per Mille.
+The modular design of VALUE also supports other multi-reward tasks, for example video recommendation in which click-through rate and watch-time completion must be balanced within a fixed generation space and dynamically weighted objectives.